# Dial-In LLM: Human-Aligned LLM-in-the-loop Intent Clustering for Customer Service Dialogues

### v3

@@ 4.2 Metrics
Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, scaled between 0 (no mutual information) and 1 (perfect correlation). However, researchers have claimed biased behavior in NMI Jerdee et al. (2024), necessitating further evaluation that reflects the true quality of the intent cluster from human perception.
@@ 4.3 Implementations
Four open-sourced Chinese LLMs [^2] The models are available at: Qwen2.5-7B; Qwen2.5-14B; Baichuan2-7B ; ChatGLM3-6b are fine-tuned using LoRA Hu et al. (2022) on 4 $\times$ Nvidia A100 GPUs. For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used. For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with intent labels. Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering. To reduce computational load, both random and convex sampling methods are implemented to select 20 representative sentences per cluster as input for the LLM.
@@ 5.2 Main Results
Table 2 presents the main results evaluated using the proposed dataset. Among the LLM-guided clustering baselines, the data-centric Keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al. (2024). However, deriving pairwise constraints for fine-tuning the embedding model, as in ClusterLLM, was ineffective due to the excessive number of clusters in the dataset. Interestingly, goodness evaluation does not always align with the NMI score. For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model. This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
@@
Notably, the proposed method does not require prior knowledge of the number of clusters and performs parameter searches during clustering, unlike baseline methods that rely on the true number of clusters. This adaptability underscores its strong application potential for extracting intent clusters from noisy text corpora Akama et al. (2020).
@@
Table 5: Performance (NMI) on English benchmarks.
@@ 5.4 Application Performance
One objective of intent clustering is to create a high-quality labeled dataset for training intent classifiers that can effectively handle future data inputs Gung et al. (2023). To demonstrate the practical advantages of the proposed methods, we further evaluate their performance by training BERT classifiers (bert-base-chinese model) based on the clustered results. Compared to the training data produced by the base model, which achieved only 65% classification accuracy, the proposed LLM-in-the-loop method improved the accuracy by 12%, providing promising evidence of the improved cluster quality.
@@ 5.5 Analysis of Computational Cost
We systematically compare the cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, with a test set of $S=3,080$ sentences and true cluster number $N=77$. The proposed LLM-in-the-loop iterative method (Algorithm 1) uses candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ LLM calls per iteration, and with $T=3$ iterations, totals $T\cdot\sum_{n_{i}\in N}n_{i}=480$ calls, each with 20 sentences input for coherence evaluation. ClusterLLM, employing a fixed triplet sampling approach with $Q=1,024$, incurs a constant $1,024$ LLM calls. Keyphrase expansion generates one keyphrase per sentence, resulting in a linear cost of $3,080$ LLM calls for the entire dataset. While our method requires more input tokens per LLM call, the cost remains efficient compared to its counterparts since the output token (i.e., “good/bad” label and intent label) is kept concise with fewer LLM calls.
@@ 5.6 Ablation Studies
Effective Coherence Evaluation with Data Sampling and LLM Crowdsourcing
Optimizing the input to LLM evaluator is crucial for enhancing performance and reducing costs Song et al. (2025); Jiang et al. (2024). As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences for coherence evaluation. Interestingly, as the sample time $t$ for repeated evaluation with majority voting or the dimension of the convex hull $d$ increases, performance consistently declines. This suggests that simpler, more efficient parameters lead to better accuracy. This pattern is verified through experiments with the ChatGLM model (see red line), which exhibits the same pattern.
@@
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Zhang et al. (2025). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by four fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
LLM qwen7b qwen14b baichuan2-7b chatglm3-6b
Accuracy 62.50% 65.00% 58.50% 60.00%
Table 6: Performance of fine-tuned LLMs in evaluating semantic coherence of intent clusters using numerical 1–4 scale scoring.
Binary Judgment vs. Numerical Scoring
To identify the optimal evaluation strategy for our clustering framework, we compared a numerical 1–4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment. We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset. The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting. The 1–4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human evaluation. In contrast, the simpler binary good/bad judgment provided consistent and interpretable quality assessments, proving to be the more effective evaluation strategy.
@@
Table 7 evaluates the impact of distance measures and merging decisions on the proposed cluster merging technique. Results highlight the superiority of geodesic distance in hyperspherical space, demonstrating its ability to effectively capture deeper semantic relationships compared to cosine similarity. Additionally, the probabilistic merging strategy consistently improves robustness by accounting for uncertainty, yielding better clustering quality regardless of the distance measure used. These findings emphasize the importance of both an appropriate distance metric and a probabilistic approach in achieving optimal merging outcomes.
{CJK}
UTF8gkai
Top1 Top2 Top3 Top4 Top5
@@
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Appendix 13). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates. This approach substantially reduces redundant model fittings and enhances the efficiency of deploying the proposed methods.
@@ 6 Conclusion
This paper tackles dialogue intent clustering using a human-aligned LLM-in-the-loop approach. Experiments on a complex, realistic Chinese customer service dataset reveal that fine-tuned LLM utilities excel in semantic coherence evaluation and cluster labeling, with the proposed methods consistently outperforming existing LLM-guided clustering baselines in cluster quality and computational efficiency. Future work should focus on refining evaluation strategies to enhance semantic guidance from LLMs and exploring task-centric approaches for deeper LLM-in-the-loop integration.
@@
This work has two main limitations. First, while the coherence evaluation is practical and intuitive, relying solely on cluster coherence as the quality indicator overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important in assessing a cluster’s usefulness. For instance, a cluster labeled “express-feeling” may be overly general and should be refined into more specific categories like “express-appreciation” and “state-dissatisfaction” to enhance interpretability and practical applicability. Additionally, the binary evaluation and numerical scoring considered in this paper are deterministic indicators, overlooking probabilistic judgments and confidence intervals, which could provide deeper insights into clustering assessment and offer greater flexibility for integration into the LLM-in-the-loop clustering process.
Second, the dataset proposed in this study is limited to the Chinese language and focuses on specific domains, namely banking, insurance, and telecommunication. While it represents the largest realistic customer service call dataset for the intent clustering problem, its coverage is restricted in terms of multi-domain and multilingual aspects, potentially affecting the generalizability of our findings to other languages and domains. Furthermore, the scope of this paper is confined to intent clustering, which serves as the initial step in the broader process of intent discovery. Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, are not comprehensively discussed. These areas warrant further investigation to fully reveal the practical values of LLM-in-the-loop design in intent mining tasks.
@@ Appendix A Data Collection and Annotation
This section details the three-stage data collection and annotation process for the proposed Chinese dialogue intent clustering dataset.
@@
The original dialogue dataset is sourced from audio transcriptions of customer service calls across several domains. Initially, the dataset contains over 11,879 customer service calls. Strict filtering is applied to maintain data integrity and exclude confidential or sensitive content, resulting in 8,184 dialogues and 69,839 sentences. Subsequent data cleaning removes duplicate sentences, resulting in a final set of 55,085 unique sentences.
@@
To construct the intent clustering dataset, we internally recruited 15 human experts, each with over five years of experience in customer service call centers, to participate in the data annotation process:
@@
    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” Five samples were provided as guides. After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.

### v4

@@ 4.2 Metrics
Normalized Mutual Information (NMI) measures the degree of similarity between ground-truth and predicted clusters, ranging from 0 (no mutual information) to 1 (perfect correlation). However, recent work has demonstrated that NMI exhibits biased behavior, particularly in favor of larger numbers of clusters Jerdee et al. (2024), highlighting the need for complementary evaluation metrics that better reflect human-perceived clustering quality.
@@ 4.3 Implementations
Four open-sourced Chinese LLMs [^2] The models are available at: Qwen2.5-7B; Qwen2.5-14B; Baichuan2-7B ; ChatGLM3-6b are fine-tuned using LoRA Hu et al. (2022) on 4 $\times$ Nvidia A100 GPUs. For coherence evaluation, a human-annotated training dataset of 1,772 intent clusters labeled as “good” or “bad” is used. For cluster naming, a training dataset of 2,500 clusters, each containing 20 sentences, is annotated with the “Action-Objective” labels. Semantic-rich embeddings are generated using the BAAI General Embeddings (BGE) model [^3] https://huggingface.co/BAAI/bge-large-zh-v1.5 Xiao et al. (2024). The best-performing clustering algorithm, hierarchical clustering, and LLM utilities, qwen14b and chatglm3-6b, are selected for LLM-in-the-loop clustering. Additionally, we employ random and convex sampling to select 20 representative sentences per cluster as LLM inputs, reducing computational overhead while preserving intent coverage.
@@ 5 Results and Discussions
LLM
qwen7b
qwen14b
baichuan2-7b
chatglm3-6b
Accuracy 96.25% 97.50% 89.17% 95.83%
LLM
qwen7b
qwen14b
baichuan2-7b
chatglm3-6b
Accuracy 92.8% 94.3% 94.3% 94.4%
Table 3: Performance of fine-tuned LLMs in naming intent clusters.
@@ 5.1 LLM Utilities
Method NMI (Mean $\pm$ Std) NMI Gain #Good (Mean $\pm$ Std) #Good Gain
Baselines
K-Means 0.7899 $\pm$ 0.0135 - 94.8% $\pm$ 1.3% -
GMMs 0.7903 $\pm$ 0.0140 +0.05% 91.1% $\pm$ 1.6% -3.90%
Hierarchical 0.8001 $\pm$ 0.0128 +1.29% 94.9% $\pm$ 1.2% +0.11%
LLM-Guided Clustering Baselines
ClusterLLM Zhang et al. (2023) 0.7284 $\pm$ 0.0168 -7.79% 91.2% $\pm$ 1.8% -3.80%
IDAS De Raedt et al. (2023) 0.8109 $\pm$ 0.0105 +2.66% * 93.6% $\pm$ 1.2% -1.27%
Keyphrase Viswanathan et al. (2024) 0.8371 $\pm$ 0.0098 +5.97% *** 94.5% $\pm$ 1.0% -0.32%
LLMEdgeRefine Feng et al. (2024) 0.7411 $\pm$ 0.0155 -6.19% 87.2% $\pm$ 2.0% -8.02%
Proposed Method (Context-Free)
LLM-ITL (random) 0.8202 $\pm$ 0.0095 +3.84% ** 97.7% $\pm$ 0.6% +3.06% ***
LLM-ITL (convex) 0.8208 $\pm$ 0.0090 +3.92% ** 97.8% $\pm$ 0.5% +3.16% ***
LLM-ITL + keyphrase 0.8378 $\pm$ 0.0085 +6.06% *** 96.4% $\pm$ 0.6% +1.69% *
Proposed Method (Context-Aware)
LLM-ITL + merge 0.8420 $\pm$ 0.0178 +6.59% ** 97.8% $\pm$ 1.3% +3.16% ***
LLM-ITL + role 0.8679 $\pm$ 0.0068 +9.86% *** 97.2% $\pm$ 0.5% +2.53% **
LLM-ITL + role + merge 0.8826 $\pm$ 0.0060 +11.76% *** 97.6% $\pm$ 0.4% +2.95% ***
Table 4: Performance of baselines and proposed methods in dialogue intent clustering. Gains are computed relative to the K-Means baseline. The best results in each column are bolded.
@@ 5.2 Main Results
Table 4 reports the main results on the proposed dataset, with all evaluation metrics averaged over five random seeds. Among the LLM-guided clustering baselines, the data-centric keyphrase expansion method, which refines sentence embeddings via LLM-summarized keyphrases, achieved the most significant performance gain Viswanathan et al. (2024). However, deriving pairwise constraints for fine-tuning the embedding model, as in ClusterLLM, was ineffective due to the excessive number of clusters in the dataset. Interestingly, goodness evaluation does not always align with the NMI score. For example, while keyphrase expansion improved NMI, the generated keyphrases could distort the original sentence’s meaning, leading to less representative clusters and a slightly lower goodness score compared to the base model. This highlights the need to balance ground truth alignment with semantic coherence during evaluation.
@@
Notably, the proposed method does not require prior knowledge of the number of clusters and automatically performs parameter searches during clustering, unlike baselines that depend on the true cluster count. This adaptability highlights its strong application potential for extracting intent clusters from noisy text corpora Akama et al. (2020).
@@
Table 5: NMI (%) performance of different clustering methods on four English intent benchmarks.
@@ 5.3 Evaluation on English Benchmarks
Method Accuracy Performance Gain
K-Means (Baseline) 0.65 -
ClusterLLM 0.62 -4.62%
LLMEdgeRefine 0.63 -3.08%
IDAS 0.68 +4.62% *
Keyphrase Expansion 0.72 +10.77% ***
LLM-ITL (context-free) 0.73 +12.31% ***
LLM-ITL (context-aware) 0.77 +18.46% ***
Table 6: Accuracy of BERT classifiers trained on datasets generated by different clustering methods.
@@ 5.4 Application Performance
One objective of intent clustering is to build a high-quality labeled dataset for training intent classifiers that can handle future inputs Gung et al. (2023). To evaluate the practical effectiveness of our methods, we trained BERT classifiers (bert-base-chinese) on clustered data generated by different approaches.
As shown in Table 6, the baseline method only achieved 65% accuracy. ClusterLLM and LLMEdgeRefine performed slightly worse, likely due to their heavier reliance on LLM-driven modifications, which can introduce label noise or overly fine-grained distinctions that reduce cluster consistency. In contrast, IDAS applies a more conservative refinement strategy, leading to a modest but stable improvement. Our proposed LLM-in-the-loop methods substantially outperformed existing approaches: LLM-ITL reached 73% accuracy, and its enhanced variant achieved 77%, representing an 18.46% relative improvement. These results are consistent with the observed cluster quality and provide additional empirical evidence that our method produces high-quality, human-aligned intent clusters with significant practical advantages.
@@ 5.5 Analysis of Computational Cost
We compare the computational cost of the proposed method with existing LLM-guided clustering on the Bank77 dataset, which contains $S=3,080$ test sentences and $N=77$ true clusters. Our LLM-in-the-loop approach evaluates candidate cluster numbers $N=[10,30,50,70]$, requiring $\sum_{n_{i}\in N}n_{i}=160$ calls per iteration. With $T=3$ iterations, this totals 480 calls, each processing 20 sentences for coherence evaluation. Cluster naming requires additional LLM calls based on the final cluster count, resulting in approximately 560 calls. In contrast, ClusterLLM uses a fixed triplet sampling strategy with $Q=1,024$, resulting in a constant cost of 1,024 calls. Keyphrase expansion generates one keyphrase per sentence, totaling 3,080 calls. While our method processes more input tokens per call, it remains cost-efficient as the output is limited to a short “good” or “bad” label or an intent label.
@@ 5.6 Ablation Studies
Data Sampling and LLM Crowdsourcing for Effective Coherence Evaluation
Optimizing input to the LLM evaluator is crucial for improving performance and reducing costs Song et al. (2025); Jiang et al. (2024). In this ablation study, we compare various sampling techniques to identify best practices for coherence evaluation. As shown in Figure 3, convex sampling outperforms random sampling in selecting representative sentences. Notably, increasing either the sample repetitions $t$ for repeated validation or the convex hull dimension $d$ consistently reduces performance. This suggests that simpler parameter choices yield higher accuracy, a trend also seen with the ChatGLM model (red line) and summarized in Table 15 with numerical comparison.
@@
To better assess the consistency and alignment between LLM judgments and human perceptions, we utilize the concept of crowdsourcing for collaborative coherence evaluation Hong et al. (2025b). The proposed LLM crowd consists of diverse entities Zhang et al. (2024) represented by four fine-tuned LLMs and uses a majority voting mechanism to aggregate their judgments Schoenegger et al. (2024). Based on the evaluation, we argue that a single fine-tuned LLM can effectively serve as a robust evaluator, as evidenced by the comparable performance to crowdsourced accuracy (cross markers in Figure 3) when using convex sampling.
@@
Table 7 evaluates the effects of distance measures and merging strategies on the proposed cluster merging method. The results show that geodesic distance in hyperspherical space outperforms cosine similarity by capturing deeper semantic relationships. Furthermore, the probabilistic merging strategy consistently enhances robustness by accounting for uncertainty, leading to better clustering quality with any distance measure. These findings underscore the value of combining an appropriate distance metric with a probabilistic approach to achieve optimal merging performance.
Top1
Top2
Top3
Top4
Top5
@@
In practice, the search space for the number of clusters can be extensive, ranging from a few options to several hundred. To mitigate the computational costs associated with traditional model-based selection, we propose an LLM-native method for search space pruning, which predicts the optimal cluster number for subsequent iterations using logs from previous iterations (see Table 16). By pruning the search space and selecting the top 5 non-repetitive predicted solutions, we achieved an accuracy of 89.47% (see Table 8), indicating a high likelihood that the optimal cluster number is among predicted candidates. This approximation approach significantly reduces redundant model fittings and improves the efficiency of applying the proposed methods to large-scale datasets.
Binary Judgment vs. Numerical Scoring for Coherence Annotation
Finally, to identify the optimal annotation strategy for coherence evaluation, we compared a numerical 1 – 4 scoring system (1: very poor, 4: very coherent) with the proposed binary good/bad judgment. We annotated 1,000 clusters with input from five human experts, finalized scores through majority voting, and fine-tuned four Chinese LLMs using this dataset. The models were then evaluated on 200 additional clusters, assigning scores five times per cluster and consolidating results via majority voting (see results in Table 9). The 1 – 4 scale demonstrated low accuracy and inconsistency, with 19% of samples observed to have significant disagreement (e.g., three ‘1’s and two ‘4’s) in Qwen-2.5-14B, making it unreliable for assessing cluster quality or aligning with human judgment. In contrast, the simpler binary good/bad judgment provided more consistent, interpretable, and reliable quality assessments, demonstrating its superiority as an evaluation protocol.
LLM
qwen7b
qwen14b
baichuan2-7b
chatglm3-6b
Accuracy 62.50% 65.00% 58.50% 60.00%
Table 9: Performance of fine-tuned LLMs in assessing cluster coherence using a numerical 1 – 4 scale.
To further validate the effectiveness and robustness of our annotation protocol, we conducted an inter-annotator agreement study with five experts. They annotated 100 intent clusters generated by the K-means algorithm, covering 2,000 sentences from recent customer service dialogues. Each cluster was evaluated using (i) a binary Good/Bad judgment and (ii) a finer-grained 1 – 4 rubric. The Fleiss’ kappa ($\kappa$) for the binary scheme reached 0.82, indicating “almost perfect” agreement Landis and Koch (1977), while the 1 – 4 rubric achieved only 0.59 (“moderate” agreement) due to variability in intermediate scores (2 and 3). These results show that binary annotation provides a more reliable and interpretable measure of cluster coherence, reinforcing the robustness of our methodology.
@@ 6 Conclusion
This paper tackles dialogue intent clustering through a human-aligned LLM-in-the-loop framework. Experiments on a large-scale Chinese customer service dataset demonstrate that fine-tuned LLM utilities are highly effective for semantic coherence evaluation and cluster labeling, enabling consistent improvements over existing LLM-guided baselines in both clustering quality and computational efficiency. Beyond achieving state-of-the-art performance, our study offers strong empirical evidence for the effectiveness of LLM-in-the-loop methodologies, with ablation studies highlighting best practices. Future work should refine evaluation beyond coherence to capture interpretability and expressiveness, extend the clustering framework to multilingual settings, and explore deeper task-centric integration of LLMs to further advance intent mining in real-world applications.
@@
Despite the effectiveness of the proposed LLM-in-the-loop intent clustering method, this study has several limitations. First, while cluster coherence is a practical and intuitive quality indicator, its sole reliance overlooks other critical attributes, such as meaningfulness and expressiveness, which are equally important for assessing the quality of intent clusters. For example, a cluster labeled “express-feeling” may be too broad and could be refined into more specific clusters like “express-appreciation” to improve interpretability and applicability. Additionally, the binary evaluation and numerical scoring used are both deterministic metrics with fixed scales, neglecting probabilistic judgments and confidence intervals that could provide deeper assessment insights and enhance flexibility for LLM-in-the-loop integration.
Second, the proposed intent clustering dataset is limited to the Chinese language and specific domains. Although it excels in capturing large-volume, realistic customer service dialogues, its scope restricts multi-domain and multilingual generalizability, potentially limiting the applicability of findings to other languages and domains. Moreover, this paper focuses solely on intent clustering, an initial step in the broader intent discovery process. Subsequent steps, such as analyzing intent trajectories and recognizing intents at the document level rather than the sentence level, encourage further investigation with more diverse LLM utilities and varied integration strategies to better reveal the practical value of LLM-in-the-loop methodologies.
Acknowledgments
This paper is partially supported by several ongoing projects led or coordinated by Prof. Zhang Chen, including P0045948 and P0046453 (industry donations from Accel Group Holding Limited and Minshang Creative Technology Holdings Limited), P0046701 and P0046703 (PolyU internal research funding), P0048183 and P0048191 (Research Matching Grant Scheme funded by the University Grants Committee), P0048887 (Innovation and Technology Fund - ITSP, ITS/028/22FP), P0051906 (RGC Early Career Scheme, 25600624), and P0054482 (industry donations from Two Square Capital Limited).
@@ Appendix A Data Collection and Annotation
This section details the three-stage data collection and annotation process for the proposed Chinese customer service dialogue intent clustering dataset.
@@
The raw dialogue dataset was derived from audio transcriptions of customer service calls in three major domains: banking, telecommunication, and insurance. In total, the dataset contained 11,879 calls. To ensure data integrity and protect confidentiality, strict filtering was applied to exclude sensitive content, resulting in 8,184 dialogues with 69,839 sentences. Further cleaning was conducted to remove duplicate sentences, yielding a final set of 55,085 unique sentences.
@@
To construct the intent clustering dataset, we recruited 15 human experts to conduct the annotations. Each annotator has over five years of professional experience in customer service call centers, with domain expertise aligned with the dataset domains: banking, telecommunication, and insurance. All annotators are fluent in Chinese, familiar with domain-specific terminology, and experienced in dialogue annotation or quality assurance tasks. Prior to annotation, a mandatory training session was held, during which the authors provided 50 intent clusters annotated with Good/Bad judgments and Action–Objective intent labels as demonstration and reference examples. This training ensured a consistent understanding of the guidelines across all annotators. The annotation process is outlined as follows:
@@
    For the good clusters, annotators were asked to label the underlying intent using the naming convention “Action-Objective.” After intent annotation, 1255 clusters with unique intentions were retained, and 28 clusters were merged due to replicated intent.